Suppose X, Y are independent and X~N(1,4) and Y N(1,9). If P(2X Y a) P(4X - 2Y 2 4a), then find a
Answer #1

Here,

E( 2X + Y) = 2E(X) + E(y) = (2 * 1) + 1 3

and Var(2X + Y) = 22V,ar(X) + Var(Y) = (22 * 4) + 9 = 25

So, 2X + Y ~ Normal( 3, 52)

Also,

E(4X-2Y) = 4E(X) - 2E(Y) = (4*1) - (2*1) = 2

and Var(4X-2Y) = 42Var(X)+22Var(Y) (42*4)+(22+9) = 64+36 = 100

So, 4X - 2Y ~ Normal(2, 102)

Now,

P(2X+ Y leq a) = P(4X - 2Y geq 4a)

Rightarrow P(rac{2X+ Y - 3}{5} leq rac{a -3}{5} ) = P(rac{4X - 2Y - 2}{10} geq rac{4a - 2}{10})

Rightarrow P(Z leq rac{a -3}{5} ) = P(Z geq rac{4a - 2}{10})

Rightarrow P(Z leq rac{a -3}{5} ) = 1 - P(Z leq rac{4a - 2}{10})

Rightarrow Phi(rac{a -3}{5} ) = 1 - Phi(rac{4a - 2}{10})

Rightarrow Phi(rac{a -3}{5} ) = 1 - Phi(rac{2a - 1}{5})

Rightarrow Phi(rac{a -3}{5} ) = Phi(-rac{2a - 1}{5})

Rightarrow rac{a -3}{5} = -rac{2a - 1}{5}

Rightarrow (a-3) = -(2a-1)

Rightarrow a-3 = -2a + 1

Rightarrow 3a= 4

Rightarrow a= rac{4}{3} (ans)